OpenStudy (anonymous):

Will give Medal! Directions: Drag each tile to the correct box. In hoping to make it onto the school's Gold Team, Bryan, Eli,and Damien took golf lessons over the summer. Bryan signed up to take 4 lessons with a semi-pro instructor, and each lesson cost $32.00 for a 45-minute session. The instructor advised Bryan to buy a bag of bulk practice balls to use during his lessons. After buying a bag of 32 golf balls, the total cost for his lessons and golf balls is $132.48. Eli signed up to take 6 lessons with a pro instructor, and each lesson cost $52.00 for a 1-hour session. Eli bought a bag of bulk practice balls to use during his lessons. After buying a bag of 40 golf balls, the total cost for his lessons and golf balls is $319.20. Damien signed up to take 5 lessons with pro instructor, and each lesson cost $28.00 for a 30-minute session. Based on the instructor's advise, he bought a bag of 24 golf balls to use during his lessons. The total cost for his lessons and golf balls is $144.08. Place the students in order from the least amount to the greatest amount that they paid per golf ball.

OpenStudy (danjs):

Lets take this one person at a time.

OpenStudy (anonymous):

Ok

OpenStudy (danjs):

Brian: 4 lessons 32 dollars per lesson Bought 32 Balls, at X dollars each Total money spent = 132.48

OpenStudy (anonymous):

Ok

OpenStudy (danjs):

In general: Total Spent = Cost of lesson*(# lessons) + (cost of golf ball X)*(#golf balls)

OpenStudy (anonymous):

Got it...

OpenStudy (danjs):

Brian: \[\large\color{blue}{ $132.48 = 4 * 32 + 32*X}\]

OpenStudy (danjs):

X= cost of 1 golf ball

OpenStudy (anonymous):

Ok

OpenStudy (danjs):

I get \[\large\color{blue}{ Brian: 32~ balls~~~@ ~~x=0.14 ~cents~each}\]

OpenStudy (anonymous):

Ok but is that the answer or are we going through all?

OpenStudy (danjs):

Have to do this for each person... so far we figured Brian payed X= $0.14 per golf ball

OpenStudy (danjs):

u see how all that was figured out though?

OpenStudy (anonymous):

Oh ok... LOL!

OpenStudy (danjs):

Moving on to Elli: Just write down the info, read one sentence at a time.... 6 Lessons $52 per lesson 40 golf balls @ $x per ball Total Spent = $319.20

OpenStudy (anonymous):

Ok

OpenStudy (danjs):

Apply the same logic to form an equation... Total Spent = (#lessons)*(Price per lesson) + (# golf Balls)*(Price Per Ball)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Okey dokey

OpenStudy (danjs):

ELI 319.20 = 6*52 + 40*x

OpenStudy (danjs):

can you solve that for x = (price per ball)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Is it 319.98?

OpenStudy (danjs):

\[319.20 = 6*52 + 40x\]\[319.20 = 312 + 40x\]\[319.2 - 312 = 40x\]\[7.2 = 40x\]\[x = \frac{ 7.2 }{ 40 } = 0.18\]

OpenStudy (anonymous):

Oh...

OpenStudy (danjs):

so ELI payed 18 cents per ball

OpenStudy (danjs):

Damien: 5 Lessons $28.00 per lesson 24 golf balls @ $x per ball total spent = $144.08

OpenStudy (danjs):

For Damien 144.08 = 5*28 + 24x

OpenStudy (danjs):

x = 0.17 cents per ball

OpenStudy (anonymous):

Sorry my test logged me out. I have to restart and wont give me the same question. :-0 But the new question I need help in too.

OpenStudy (danjs):

OVERALL: Brian = $0.14 per ball Eli = $0.18 per ball Damien = $0.17 per ball

OpenStudy (danjs):

oh, it will prolly be similar though wont it

OpenStudy (anonymous):

Yes it is kinda like it but not too too much. LoL!

OpenStudy (anonymous):

Here it is: At their going-out-of-business sale, Bill's Bikes sold 36 bikes and had 54 left. Which equation could be used to find x, the number of bikes the store had originally? x - 54 = 36 x + 36 = 54 x + 54 = -36 x - 36 = -54

OpenStudy (danjs):

ok

OpenStudy (danjs):

Sold 36 now has 54 left some number - 36 = 54 left

OpenStudy (danjs):

so the second choice

OpenStudy (anonymous):

Thank you! I have another one.

OpenStudy (anonymous):

Cormac is making 29 phone calls each week as part of a fundraiser. Last week, he made 11 phone calls on weekdays. Which equation could be used to find x, the number of phone calls he made last week on weekend days? 29x = 11 x + 29 = 11 11 + x = 29 x - 29 = 11